Exercise 12-7 In early January 2016, Bramble Corporation applied for a trade name, incurring legal costs of $16,080. In January 2017, Bramble incurred $7,880 of legal fees in a successful defense of its trade name Compute 2016 amortization, 12/31/16 book value, 2017 amortization, and 12/31/17 book value if the company amortizes the trade name over 10 years. (Round all answers to 0 decimal places, e.g. 8,564) 2016 amortization 12/31/16 book valuel 2017 amortization 12/31/17 book vaiue s

Compute the 2017 amortization and the 12/31/17 book value, assuming that at the beginning of 2017, Bramble determines that the trade name will provide no future benefits beyond December 31, 2020. 2017 amortization 12/31/17 book value LINK TO TEXTLINK TO TEXT VIDEO: SIMILAR EXERCISE Ignoring the response for part (b), compute the 2018 amortization and the 12/31/18 book value, assuming that at the beginning of 2018, based on new market research, Bramble determines that the fair value of the trade name is $14,560. Estimated total future cash flows from the trade name is $15,760 on January 3, 2018. 2018 amortization 12/31/18 book value
Answer #1
1. 2016 Amortization = $16080/10years
= $1608
12/31/2016 Book Value = $16080-$1608 = $14472
2017 Amortization = ($14472 +$7880)/9 = $2483.56
13/31/2017 Book Value = ($14472+$7880-2483.56) = $19868.44
B) Assuming trade names will provide no future benefits beyond December 2020
2017 Amortization = ($14472 +$7880)/4 = $5588
12/31/2017 Book Value = ($14472 +$7880- $5588)= $16764
C)
Carrying Amount 19868.44 is greater than future cash flows $15760, thus the Trade name fails the recovery test. The new carrying value is $14560, the trade name's fair value
2018 Amortization (After recording impairment loss)
$14560/8 = $1820
12/31/2018 Book Value = ($14560-$1820) = $12740